# Searching for GeV Gamma-Ray Polarization and Axion-Like Particles with AMS-02

### v1

@@ IV.1 Sensitivity to polarization fraction
where $A=0.14$ for AMS-02 as defined before, $N_{S}$ and $N_{B}$ are signal and background counts respectively, and $p$ is the probability threshold for a detection, i.e. above the MDP($p$) polarization threshold, the associated modulation amplitude has probability $\leq p$ that it would be exceeded by a statistical fluctuation if the true polarization fraction were zero. Note that while the MDP is commonly used as an estimate of the sensitivity of a polarimeter, it is not the same as (for example) “the true polarization value for which the zero-polarization hypothesis is expected to be excluded at the $1-p$ confidence level” (see also [54] for discussion); the latter quantity depends on the uncertainty in the extraction of the polarization amplitude when the true polarization is non-zero, whereas MDP is calculated purely under the zero-polarization hypothesis. However, it is useful as an estimate for the number of counts needed to render polarization potentially detectable.
@@
As an optimistic estimate, we take $N_{B}=0$, since all the sources we consider are bright relative to the local diffuse background and the background from charged cosmic ray is at a level lower than the local diffuse background [24]. We see that for the MDP to be below 1 with $p=0.05$, we need $N_{S}>(2\sqrt{-\ln 0.05}/0.14)^{2}\approx 600$. This sets a floor for the brightness of sources where we can hope to measure polarization.
For each of the sources we consider, we compute the expected number of photons in 10 years of AMS-02 data. We estimate the total detected events by integrating the emission spectra multiplied by the acceptance of the AMS-02 detector (described in Ref. [24]) over a period of ten years (corresponding to the operation time of the AMS-02 detector). As for the AMS-100 detector, we assume its acceptance is 100 times of that of AMS-02. In Table 4 we show the resulting MDP for each of our selected sources using AMS-02. We should note that for some of the sources, AMS-02 will not have enough sensitivity to distinguish an unpolarized state even from a fully polarized state. We see that for both the highest-statistics pulsars and supernova remnants, the MDP is around $30\%$; i.e. a polarization fraction above $\sim 30\%$ would correspond to an amplitude that has less than $5\%$ probability if the emission is unpolarized. For example, for the supernova remnant W44 we expect 5542.31 events in 10 years of AMS-02 data, corresponding to a MDP of 0.33 at 95% confidence.
@@
Table 4: Minimal detectable polarization for selected sources. The first six entries correspond to pulsars, the next four to supernovae, and the last to our example extragalactic source. 1 should be understood as there is no way to distinguish a polarized source from an unpolarized one with the statistics.
Source name Estimated counts for AMS-02 in ten years minimal detectable polarization for AMS-02 minimal detectable polarization for AMS-100
J1420-6048 1726.45 0.60 0.06
J1648-4611 317.35 1 0.14
J1702-4128 369.33 1 0.13
J1718-3825 1175.91 0.72 0.07
J2021+3651 5209.57 0.34 0.03
J2240+5832 294.85 1 0.14
IC 443 4376.49 0.37 0.04
W44 5542.31 0.33 0.03
W51C 3085.58 0.45 0.04
W49B 1360.96 0.67 0.07
NGC1275 2775.22 0.47 0.05
@@ IV.2 Monte Carlo simulations for ALP parameter space
For each of the sources discussed in this work, and for each point $\{m_{a},g_{a\gamma\gamma}\}$ in ALP parameter space, we calculate the expected photon polarization spectrum at the AMS-02 detector using the modeling described in section II. We compute the expected counts and their polarization fraction over the full 0.1-10 GeV energy range, and in four or ten log-spaced energy bins, using the detector properties described in section III. For each choice of binning, we average the forecast linear polarization fractions as a scalar sum over the total counts within each bin (weighted by the number of photons as a function of energy) and use it as the true polarization fraction for the Monte Carlo (MC) simulations. To treat it more rigorously, one should average the individual polarization+intensity vectors, using a vector sum over counts instead. In practice, we have confirmed that since the polarization angle varies continuously without sudden jumps, the two averaging methods yield similar results. The spectral distortion due to the bin-to-bin migration is negligible for these wide bins.
@@
Fig. 5 shows an example of a simulated event distribution (in azimuthal angle) for a unbinned (in energy) analysis for the W44 supernova remnant, assuming initially unpolarized photons with ALP parameters $g_{a\gamma\gamma}=3.543\times 10^{-10}$ GeV-1 and $m_{a}=4.41\times 10^{-9}$ eV. As previously, W44 gives rise to 5542.31 expected counts for the AMS-02 case, and 100$\times$ that number (554231) for ten years with AMS-100.
@@ IV.3 Statistical tests for polarization and energy dependence
  • In the case of a single large energy bin, the $p$-value of the (mock) data with respect to the zero-polarization hypothesis (which translates into the significance with which we can reject this hypothesis).
@@
To compute the $p$-values we employ a $\chi^{2}$ test. For Galactic sources, having determined these results for each realization, we average the $p$-value across the 1000 realizations to obtain the expected confidence level (for exclusion of the null hypothesis / detection of a signal), which is $1-$ (expected $p$-value).
For NGC1275 analyses, we need to take into account the uncertainties in the $B$-field model. For each ALP parameter point and each of the 100 choices of $B$-field configuration, we evaluate the expected $p$-value as described above (averaging over 1000 realizations), and then rank the $B$-field configurations by the expected $p$-value (independently for each ALP parameter point). We will show results for the median $B$-field configuration by this metric (i.e. $50\%$ of B-field configurations would predict a higher significance and $50\%$ lower), as well as for a “pessimistic” $B$-field configuration at the $95$th percentile of $p$-value (i.e. $95\%$ of B-field configurations would predict a lower $p$-value / higher significance of detection).
@@ IV.3.1 Detecting ALP-induced polarization from an initially unpolarized source
In Fig. 6 we show the result of these two calculations as a function of the true polarization, for an example source. We found that at $p=0.05$ (and lower values), the two calculations coincided closely, suggesting that the second (simpler) approach is adequate for computing the region where we expect to exclude the null hypothesis at $95\%$ confidence or better. Accordingly, we use the second method for the rest of our analyses. We attribute this agreement to the fact that exclusion of the null hypothesis at $95\%$ confidence requires both a large number of photon counts (meaning the Poisson distribution can be well-approximated by a Gaussian) and that the true polarization is well away from the boundary value of zero polarization. In this case we also find quite close agreement between the polarization degree corresponding to $p=0.05$ and the analytic MDP calculation for $p=0.05$ (Eq. 17).
It is already apparent from Fig. 6 (and from the analytic estimates of Table 4) that detection of energy dependence in the polarization fraction, from steady-state Galactic sources, will be very challenging with AMS-02. The simplest version of such an analysis would involve splitting the dataset into two bins; if we chose these bins to have roughly equal numbers of photon counts, the MDP at 95% confidence would rise to roughly 0.5 in each bin (from 0.3 in the full dataset). In the maximally optimistic case where one bin had zero polarization and the other had 100%, we would expect to only be able to say (at 95% confidence) that one bin had below 50% polarization and the other had above 50%, which is only a very marginal detection of variation. Considering realistic scenarios where the polarization varies smoothly with energy, even before specifying the ALP model, we see that AMS-02 is unlikely to have the statistical power to distinguish such scenarios.
In contrast, for AMS-100, there will be ample events for both of these analysis. As a result, we will only show projections of detection of a non-zero polarized fraction for the AMS-02 detector; whereas for the future AMS-100 detector we will include the results of both this analysis and a search for energy-dependence in the polarization signal.
[…]
@@ V.1 Extragalactic sources
Figure 8: Expected sensitivity with ten years data from NGC1275. We randomly sampled 100 $B$-field configurations and calculated the expected $p$-value for each of them; at each point in ALP parameter space we rank the B-field configurations by their corresponding expected $p$-value. The upper panels show the contours of expected confidence level (1 - expected $p$-value) for detecting ALP-induced polarization (i.e. excluding the no-polarization hypothesis), with ten years of data, for the median B-field scenario at each point, for (left panel) AMS-02 and (right panel) AMS-100. In the left panel we also show (dotted line) the contour of expected confidence level $0.95$ for the 5th best B-field scenario (of 100) for each point, corresponding to an optimistic $B$-field model. The bottom left panel shows the contours of expected confidence level for detecting polarization in the 95th best B-field scenario (of 100) at each point, for ten years of AMS-100 data, corresponding to a pessimistic $B$-field model. The bottom right panel shows the contours of expected confidence level for exclusion of the constant-polarization hypothesis, with ten years of AMS-100 data, for the median B-field scenario. The cross-hatched region shows the contour corresponding to a $p$-value of 0.05 for the 95th best B-field scenario (out of 100) at each point.
@@
We summarize the regions of ALP parameter space where we expect to be able to exclude the relevant null hypotheses, using NGC1275 observations, in Fig. 8. We show contours of the expected confidence level (1 - $p$-value) for exclusion of an unpolarized signal in AMS-02 (upper left) and AMS-100 (upper right and lower left), and for exclusion of an energy-independent polarization with AMS-100 (lower right). In the upper panels, we show the results for the median realization (ranked by $p$-value at each point) of the NGC1275 $B$-field (and additionally the result for an optimistic $B$-field configuration in the upper left panel). In the lower left panel, we show the analogous results when instead at each ALP parameter point we choose the $B$-field configuration that is ranked 95th (out of 100) in terms of the expected significance of detection, for AMS-100. Following this procedure for AMS-02 would mean the expected $p$-value is always above 0.05, i.e. there is no point in ALP parameter space where we expect AMS-02 data to exclude the no-polarization hypothesis at $95\%$ confidence or better in at least $95\%$ of possible $B$-field configurations. In the lower right panel, for the energy-dependence analysis with AMS-100, we show the region where the expected confidence level for detection exceeds 0.95 for the “pessimistic” (95/100) $B$-field configuration as a cross-hatched overlay on the contours for the median $B$-field realization. We observe that in this case, the effect of the more pessimistic $B$-field model is quite modest, in contrast with AMS-02. Comparing the two lower panels indicates that in the parameter region where we potentially have sensitivity to energy-dependence with AMS-100, high confidence levels for detection are ubiquitous even for quite pessimistic $B$-field models.
@@
Figure 9: Expected ($95\%$ confidence level) sensitivity region for signals from NGC1275 with the median $B$-field configuration, compared to other existing constraints. The shaded region indicates the projected sensitivity region for detecting an energy-dependent polarization signal from AMS-100 with ten years of data. The two red dashed lines indicate the projected sensitivity region for detecting an ALP-induced polarization signal with ten years of data from AMS-02 or AMS-100, as labeled in the plot. The possible signal from Fermi-LAT [15] is shaded in light blue. The other existing constraints are plotted using AxionLimitPlotter [55].
In Fig. 9, we compare our derived sensitivity regions to existing constraints. We find that, with the existing AMS-02 dataset and for the median $B$-field realization, the possible ALP parameter space that can be probed is already in tension with other existing constraints. For an optimistic $B$-field realization, however, it might be possible to observe ALP-sourced polarization from NGC1275 for a region of parameter space that is not yet excluded in the mass range around 0.1-1 neV; for example, we see in the upper left panel of Fig. 8 that for the 5th-best $B$-field model out of the 100 we tested (ranked by expected $p$-value), AMS-02 could have sensitivity to ALP-induced polarization for $g_{a\gamma\gamma}\gtrsim 10^{-12}\text{GeV}^{-1}$ for $m_{a}\lesssim 2\times 10^{-10}$ eV, with some sensitivity up to $m_{a}\lesssim 3\times 10^{-9}$ eV. Of course, such a detection could not be taken as unambiguous evidence of ALPs unless we could place a stringent prior constraint on the degree to which the AGN emission is itself polarized. Furthermore, these more optimistic B-field models might impact constraints on ALP-induced signals from the gamma-ray intensity spectrum of NGC1275 with Fermi-LAT data (as in e.g. [28, 18, 29]). An analysis using a common set of B-field configurations (for both Fermi-LAT and AMS-02 signals) would be needed to establish whether AMS-02 could hope to see a polarization signal without the ALP explanation being simultaneously ruled out by Fermi-LAT intensity data.
@@
An AMS-02 successor with 100$\times$ acceptance would extend its reach into parameter space that is not yet tested by any experiment, and also offer the possibility of detecting energy dependence of the polarization for $10^{-10}$ eV $\lesssim m_{a}\lesssim 10^{-8}$ eV. The region in which energy dependence could potentially be detected is similar in shape and reach to, but does not fully overlap with, the region excluded by updated analyses of gamma-ray intensity variation from NGC1275 using Fermi-LAT data [18, 29], with our AMS-100 forecast extending to slightly lower coupling; this is not surprising, as intensity variations with energy will generally be accompanied by polarization variations, and vice versa. We do not overplot these constraints directly in Fig. 9 because the analyses employ different magnetic field models, and this may drive artificial differences between the apparent constraints/sensitivity. A detailed comparison would require use of a shared set of magnetic field configurations (and would also depend on the details of the successor experiment).
@@
At lower masses and couplings there is a region of parameter space where there is no detectable energy dependence, but an ALP-induced polarization signal could be measured (by AMS-100 or a similar experiment). Here there is no competing constraint from gamma-ray intensity measurements, since for intensity measurements the signal of interest intrinsically involves energy-dependent variation.
For the polarization-detection analysis, we also tested the sensitivity of a successor experiment with only 10$\times$ the acceptance (such as ALADInO [27]), which might be more feasible than the full AMS-100 proposal. The sensitivity to $g_{a\gamma\gamma}$ at small $m_{a}$ is approximately the geometric mean of the sensitivities for AMS-02 and AMS-100, but we find that as $m_{a}$ is increased, the cutoff in sensitivity occurs at a similar $m_{a}$ value to our AMS-100 forecast. Such an experiment would also probe currently unconstrained ALP parameter space for the median $B$-field configuration.
@@
However, this argument neglects the systematic uncertainties associated with the modeling of the magnetic field. In particular, for AMS-02, a null result in a polarization search targeting NGC1275 will not imply any robust exclusion of ALP parameter space. This follows from our result that for AMS-02, for every point in ALP parameter space, in at least $5\%$ of B-field scenarios the expected confidence level for detection is below 95%. Thus we do not generally expect to be able to argue that a lack of detected polarization in AMS-02 implies the absence of ALPs (at $95\%$ confidence or better), as the absence of a signal could simply be due to a pessimistic $B$-field configuration.
The situation is more positive for AMS-100, especially in the part of parameter space with sensitivity to energy-dependence, as shown in the bottom right panel of Fig. 8. As discussed above, the presence of a signal in this region is relatively robust to uncertainties in the $B$-field modeling, and accordingly a null result (either in the overall polarization search or the search for energy-dependence) could provide constraints on ALPs. The lower-mass region of ALP parameter space where we would expect to detect polarization but no energy-dependence will be more challenging to either distinguish from astrophysics (in the presence of a polarization detection) or to firmly exclude (in the event of a null result), given the dependence on the $B$-field configuration shown in the bottom left panel of Fig. 8.

### v2

@@ IV.1 Sensitivity to polarization fraction
where $A=0.14$ for AMS-02 as defined before, $N_{S}$ and $N_{B}$ are signal and background counts respectively, and $p$ is the probability threshold for a detection, i.e. above the MDP($p$) polarization threshold, the associated modulation amplitude has probability $\leq p$ that it would be exceeded by a statistical fluctuation if the true polarization fraction were zero. Note that while the MDP is commonly used as an estimate of the sensitivity of a polarimeter, it is not the same as (for example) “the true polarization value for which the zero-polarization hypothesis is expected to be excluded with a given $p$-value” (see also [58] for discussion); the latter quantity depends on the uncertainty in the extraction of the polarization amplitude when the true polarization is non-zero, whereas MDP is calculated purely under the zero-polarization hypothesis. However, it is useful as an estimate for the number of counts needed to render polarization potentially detectable.
@@
As a first optimistic estimate, we take $N_{B}=0$. We see that for the MDP to be below 1 with $p=0.05$, we need $N_{S}>(2\sqrt{-\ln 0.05}/0.14)^{2}\approx 600$. This sets a floor for the brightness of sources where we can hope to measure polarization.
For each of the sources we consider, we compute the expected number of photons in 10 years of AMS-02 or AMS-100 data, estimating the exposure as described in the previous section. In Table 4 we show the estimated counts for each source in AMS-02. We see that of the sources we consider, only the Geminga and Vela pulsars are expected to have enough photons to be above the (zero-background) threshold at which MDP <1<1. Ignoring the background counts (which as we will show are quite small for both Geminga and Vela), the MDP is 0.52 for Geminga, and 0.53 for Vela.
What of the backgrounds? A full analysis including the diffuse photon background and the background from misidentified cosmic rays would take into account the probability that each event is associated with the source vs being part of the background, based on its distance from the center of the source and the point spread function of the instrument.
Such an in-depth analysis is beyond the scope of this article; instead, we estimate the background by computing the background within a $1^{\circ}$ radius of each source, which corresponds roughly to the 68% containment region of the AMS-02 point spread function at an energy of 700 MeV [27]. The angular resolution improves at higher energies, to slightly below 0.1 (in 68% containment angle) at 10 GeV, and degrades at lower energies, to around $3^{\circ}$ at 200 MeV (below this energy the effective area for conversion events becomes negligible). We could average this angular resolution function over energy, weighting by the expected photon counts; this inherits the energy dependence of the effective area, and we will also assume a photon source with a spectral index of $dN/dE\sim E^{-2}-E^{-3}$ (typical of the Galactic diffuse emission). For this range of photon spectra we obtain photon-number-averaged values for the 68% containment angle of $\sim 0.9-1.5^{\circ}$, supporting our use of $1^{\circ}$ as an estimate.
Before we go on to discuss the modeling of the diffuse photon background, let us briefly discuss misidentified cosmic rays. Ref. [27], which we use for our exposure estimates for AMS-02 (including selection cuts), suggests that in a broad region including the Galactic plane (Galactic latitude $|b|<8^{\circ}$) the cosmic-ray background is generally subdominant to the diffuse photon background. However, not all our sources are along the plane. Ref. [27] also provides an estimate for cosmic-ray background events based on the observed difference between the measured and modeled gamma ray signals; they show that the background is declination-dependent, but has a maximum value of around 20,000 events/sr at high declinations (at lower declinations this rate is suppressed by a factor of around two), in 6 years of AMS-02 data. This corresponds to 19 events in a $1^{\circ}$ radius circle in the 6-year dataset of that work, so roughly 32 events over 10 years. Thus, for any source where the minimum detectable polarization is below 1 (requiring $N_{S}>600$), we see that this background must be negligible ($\leq 5\%$) compared to the signal, with the ratio $\sqrt{N_{S}+N_{B}}/N_{S}$ thus changing only at the percent level due to the inclusion of this background. For this reason, we will ignore the cosmic-ray background for the remainder of our analysis.
In contrast, the diffuse photon background can be important, especially for relatively faint sources lying in the Galactic plane. We employ the Pass 8 (P8R3) model for the diffuse Galactic gamma-ray emission provided by the Fermi-LAT Collaboration [^3] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html to obtain the photon flux at the location of each source (as a function of energy), and then use the previously-derived exposure at the source location to convert this flux to a number of counts. The results are shown in Table 4, for the total counts integrated from 0.1 GeV to 10 GeV. We see that for sources in the Galactic plane, the backgrounds are often comparable to the counts from the sources themselves; however, for NGC1275 and the Geminga and Vela pulsars, which are further from the Galactic plane, the backgrounds are expected to be quite subdominant.
As a result, for the sources where polarization may be detectable with AMS-02 – i.e. the Geminga and Vela pulsars – we expect the zero-background MDP estimate to be quite accurate.
For the other sources, we will now move on to the prospects with AMS-100. Table 5 shows the estimated MDP both with and without background events. As we see, for the best sources in all categories (Galactic pulsars, Galactic supernovae, and the extragalactic source NGC1275), we expect to have MDP at the 1-4% level even in the presence of backgrounds, and the backgrounds do not markedly degrade the estimated MDP, although for other sources the impact of backgrounds can be substantial.
Finally, with these diffuse photon background results in hand, let us briefly revisit our neglect of the cosmic-ray background. We see that for Vela, Geminga, and NGC1275, the two backgrounds are of the same order; however, they are both very subdominant compared to the source brightness. For all other sources, the diffuse photon background dominates over the cosmic-ray background. Thus we expect our neglect of the cosmic-ray background to have no qualitative effect on our results. If AMS-100 sought to study faint sources far from the Galactic plane in polarization, the cosmic-ray background might then become a limiting factor.
@@
Table 5: Estimated signal and background counts for AMS-100 in ten years with minimal detectable polarization (MDP) for selected sources. The first six entries correspond to pulsars, the next four to supernovae, and the last three to our example extragalactic source and Geminga and Vela pulsars. 1 should be understood as there is no way to distinguish a polarized source from an unpolarized one with the statistics.
Source name Estimated counts for AMS-100 in ten years Estimated bkg counts for AMS-100 in ten years MDP for AMS-100 MDP for AMS-100 with bkg
J1420-6048 291744 1287312 0.05 0.11
J1648-4611 53938 1704463 0.11 0.61
J1702-4128 62565 1847690 0.10 0.55
J1718-3825 198806 1386797 0.06 0.16
J2021+3651 881567 619863 0.03 0.03
J2240+5832 49832 269155 0.11 0.28
IC443 741047 196893 0.03 0.03
W44 938385 1603193 0.03 0.04
W51C 522037 926101 0.03 0.06
W49B 230353 1174738 0.05 0.13
NGC1275 468941 44404 0.04 0.04
Geminga 8853197 49339 0.01 0.01
Vela 19042725 102915 0.01 0.01
@@ IV.2 Monte Carlo simulations for ALP parameter space
In this light, we now move on to study the degree to which AMS-100 could constrain ALP parameter space. (We do not perform this analysis for AMS-02 given the results of the simple MDP estimate, which suggest that none of the more distant sources will have the statistics required to measure any level of polarization.)
For each of the sources discussed in this work, and for each point $\{m_{a},g_{a\gamma\gamma}\}$ in ALP parameter space, we calculate the expected photon polarization spectrum at the AMS-100 detector using the modeling described in section II. We compute the expected counts and their polarization fraction over the full 0.1-10 GeV energy range, and in four or ten log-spaced energy bins, using the detector properties described in section III. For each choice of binning, we average the forecast linear polarization fractions as a scalar sum over the total counts within each bin (weighted by the number of photons as a function of energy) and use it as the true polarization fraction for the Monte Carlo (MC) simulations. To treat it more rigorously, one should average the individual polarization+intensity vectors, using a vector sum over counts instead. In practice, we have confirmed that since the polarization angle varies continuously without sudden jumps, the two averaging methods yield similar results. The spectral distortion due to the bin-to-bin migration is negligible for these wide bins.
@@
Fig. 5 shows an example of a simulated event distribution (in azimuthal angle) for a unbinned (in energy) analysis for the W44 supernova remnant, assuming initially unpolarized photons with ALP parameters $g_{a\gamma\gamma}=3.543\times 10^{-10}$ GeV-1 and $m_{a}=4.41\times 10^{-9}$ eV. As previously, W44 gives rise to 938385 expected counts over ten years with AMS-100.
@@ IV.3 Statistical tests for polarization and energy dependence
  • In the case of a single large energy bin, the $p$-value of the (mock) data with respect to the zero-polarization hypothesis.
@@
To compute the $p$-values we employ a $\chi^{2}$ test. For Galactic sources, having determined these results for each realization, we average the $p$-value across the 1000 realizations to obtain the expected $p$-value (for exclusion of the null hypothesis / detection of a signal).
For NGC1275 analyses, we need to take into account the uncertainties in the $B$-field model. For each ALP parameter point and each of the 100 choices of $B$-field configuration, we evaluate the expected $p$-value as described above (averaging over 1000 realizations), and then rank the $B$-field configurations by the expected $p$-value (independently for each ALP parameter point). We will show results for the median $B$-field configuration by this metric (i.e. $50\%$ of B-field configurations would predict a higher expected $p$-value and $50\%$ lower), as well as for an “unfavorable” $B$-field configuration at the $95$th percentile of $p$-value (i.e. $95\%$ of B-field configurations would predict a lower $p$-value / higher significance of detection).
@@ IV.3.1 Detecting ALP-induced polarization from an initially unpolarized source
In Fig. 6 we show the result of these two calculations as a function of the true polarization, for an example source. We found that the two calculations are generally in reasonable agreement, and the second (simpler) approach always slightly underestimates the constraining power of the analysis. Accordingly, we use the second method for the rest of our analyses, noting that this may lead to slightly weaker sensitivity forecasts. We attribute the fairly good agreement to the fact that achieving an expected $p$-value of 0.05 requires both a large number of photon counts (meaning the Poisson distribution can be well-approximated by a Gaussian) and that the true polarization is well away from the boundary value of zero polarization. In this case we also find quite close agreement between the polarization degree corresponding to $p=0.05$ and the analytic MDP calculation for $p=0.05$ (Eq. 18).
It is already apparent from Table 4 that detection of an ALP-induced polarization signal with AMS-02 is implausible, since polarization is only expected to be detectable at all in Geminga and Vela (which are not favorable targets for an ALP-induced polarization signal). Furthermore, the relatively high MDP even for these sources means that we do not expect to be able to obtain meaningful constraints on the energy dependence of the polarization fraction with AMS-02, even if a polarization signal is detected.
However, for AMS-100 (or AMS-100P), many sources will have sufficiently high statistics to measure both polarization and its energy dependence. As a result, from this point on we will focus solely on AMS-100/AMS-100P when forecasting sensitivity to ALP parameters.
Figure 6: The expected $p$-value for exclusion of an unpolarized signal, as a function of the photon polarization on arrival of the signal at the detector, for a photon count rate consistent with the supernova W44 (i.e. 938385 counts in AMS-100 data as blue lines and $20\%$ of that as orange lines). The dotted line corresponds to our first approach, deriving TS from the log likelihood ratio and determining the $p$-value using Monte Carlo. The solid line corresponds to the second (simplified) approach, where we make a Gaussian-likelihood approximation for the TS and approximate its distribution under the null hypothesis by a $\chi^{2}$ distribution. The dashed purple line corresponds to $p$-value=0.05.
@@ V.1 Extragalactic sources
Figure 8: Expected sensitivity with ten years data from NGC1275, for (left panels) AMS-100P and (right panels) AMS-100. We randomly sampled 100 $B$-field configurations and calculated the expected $p$-value for each of them; at each point in ALP parameter space we rank the B-field configurations by their corresponding expected $p$-value. All panels show the contours of (1 - expected $p$-value) for (upper panels) detecting ALP-induced polarization (i.e. excluding the no-polarization hypothesis), or (lower panels) detecting energy dependence (i.e. excluding the constant-polarization hypothesis), for the median B-field scenario at each point. The green line corresponds to the $p=0.05$ contour (for the median B-field scenario). The cross-hatched region in each plot shows the contour corresponding to a $p$-value of 0.05 for the 95th best B-field scenario (out of 100) at each point.
@@
We summarize the regions of ALP parameter space where we expect to be able to exclude the relevant null hypotheses, using NGC1275 observations, in Fig. 8. We show contours of the expected (1 - $p$-value) for exclusion of an unpolarized signal in AMS-100P (upper left) and AMS-100 (upper right), and for exclusion of an energy-independent polarization in the same two experimental configurations (lower left and right). In all panels, we show the results for the median realization (ranked by $p$-value at each point) of the NGC1275 $B$-field. In the cross-hatched regions we show the expected $p=0.05$ contour when instead at each ALP parameter point we choose the $B$-field configuration that is ranked 95th (out of 100) in terms of the expected $p$-value. We observe that in the analysis of energy-dependence, in particular, the effect of the more unfavorable $B$-field model is quite modest for AMS-100. In the parameter region where we potentially have sensitivity to energy-dependence with AMS-100, high confidence levels for detection are ubiquitous even for quite unfavorable $B$-field models. However, this statement is sensitive to the exact exposure: AMS-100P has almost the same sensitivity as AMS-100 in the presence of the median $B$-field model, but the sensitivity to energy-dependence is almost entirely lost for an unfavorable $B$-field model.
@@
Figure 9: Sensitivity region (expected $p\leq 0.05$) for signals from NGC1275 with the median $B$-field configuration, compared to other existing constraints. The shaded region indicates the projected sensitivity region for detecting an energy-dependent polarization signal from AMS-100 with ten years of data. The red dashed lines indicate the projected sensitivity region for detecting an ALP-induced polarization signal with ten years of data from AMS-100, as labeled in the plot. The possible signal from Fermi-LAT [18] is shaded in light blue. The other existing constraints are plotted using AxionLimitPlotter [59].
@@
In Fig. 9, we compare our derived sensitivity regions to existing constraints. We see that AMS-100, or even a smaller pathfinder such as AMS-100P, would extend its reach into parameter space that is not yet tested by any experiment, and also offer the possibility of detecting energy dependence of the polarization for $10^{-10}$ eV $\lesssim m_{a}\lesssim 10^{-8}$ eV. The region in which energy dependence could potentially be detected is similar in shape and reach to (but does not perfectly coincide with) the region excluded by updated analyses of gamma-ray intensity variation from NGC1275 using Fermi-LAT data [21, 32], with our forecasts for both AMS-100P and AMS-100 extending to slightly lower coupling; this is not surprising, as intensity variations with energy will generally be accompanied by polarization variations, and vice versa. We do not overplot these constraints directly in Fig. 9 because the analyses employ different magnetic field models, and this may drive artificial differences between the apparent constraints/sensitivity. A detailed comparison would require use of a shared set of magnetic field configurations.
@@
At lower masses and couplings there is a region of parameter space where there is no detectable energy dependence, but an ALP-induced polarization signal could be measured (by AMS-100, its pathfinder, or a similar experiment). Here there is no competing constraint from gamma-ray intensity measurements, since for intensity measurements the signal of interest intrinsically involves energy-dependent variation.
@@
However, this argument neglects the systematic uncertainties associated with the modeling of the magnetic field. As discussed above, in the region where AMS-100 would have sensitivity to energy-dependent polarization, the signal is relatively robust to uncertainties in the $B$-field modeling, and accordingly a null result (either in the overall polarization search or the search for energy-dependence) could provide constraints on ALPs. The lower-mass region of ALP parameter space where we would expect to detect polarization but no energy-dependence will be more challenging to either distinguish from astrophysics (in the presence of a polarization detection) or to firmly exclude (in the event of a null result), given the dependence on the $B$-field configuration shown in Fig. 8. The full version of AMS-100 would improve markedly on AMS-100P in this regard, for both energy-independent and energy-dependent analyses.